# Dual-Objective Reinforcement Learning with Novel Hamilton-Jacobi-Bellman Formulations

### v1

@@ VIII Conclusion
-A Proof Notation
@@
We let =$\mathbb{N}=\{0,1,\dots\}$ be the set of whole numbers.
We let $\mathbb{{A}}$ be the set of maps from $\mathbb{N}$ to $\mathcal{{A}}$. In other words, $\mathbb{{A}}$ is the set of sequences of actions the agent can choose. Given $\mathbf{a}_{1},\mathbf{a}_{2}\in\mathbb{{A}}$, and $\tau\in\mathbb{N}$, we let $[\mathbf{a}_{1},\mathbf{a}_{2}]_{\tau}$ be the element of $\mathbb{{A}}$ for which
$[\mathbf{a}_{1},\mathbf{a}_{2}]_{\tau}(t)=\begin{cases}\mathbf{a}_{1}(t)&t<\tau,\\ \mathbf{a}_{2}(t-\tau)&t\geq\tau.\end{cases}$
Similarly, given $a\in\mathcal{{A}}$ and $\mathbf{a}\in\mathbb{{A}}$, we let $[a,\mathbf{a}]$ be the element of $\mathbb{{A}}$ for which
$[a,\mathbf{a}](t)=\begin{cases}a&t=0,\\ \mathbf{a}(t-1)&t\geq 1.\end{cases}$
@@
The $[\cdot,\cdot]_{\tau}$ operation corresponds to concatenating two action sequences (using only the $0^{\text{th}}$ to $(\tau-1)^{\text{st}}$ elements of the first sequence), the [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] operation corresponds to prepending an action to an action sequence, and the $\cdot|_{\tau}$ operation corresponds to removing the $0^{\text{th}}$ to $(\tau-1)^{\text{st}}$ elements of an action sequence.
@@
-B Proof of RAA Main Theorem
@@
[…]
@@
We introduce some useful notation we will use throughout the rest of the proof. For each $s\in\mathcal{{S}}$, let $[s]^{+}=f(s,\pi(s))$, $[y]^{+}_{s}=\max\{y,r\left([s]^{+}\right)\}$, $[z]^{+}_{s}=\min\{z,q\left([s]^{+}\right)\}$.
@@
    In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ for $t<T$. Then $\bar{x}_{t}=x^{\circ}_{t}$, $\bar{y}_{t}=y^{\circ}_{t}$, and $\bar{z}_{t}=z^{\circ}_{t}$ for all $t\leq T$. It follows that $[\bar{x}_{t}]^{+}=x^{\circ}_{t+1}$, $[\bar{y}_{t}]_{\bar{x}_{t}}^{+}=y^{\circ}_{t+1}$, and $[\bar{z}_{t}]_{\bar{x}_{t}}^{+}=z^{\circ}_{t+1}$ for all $t\leq T$. Thus by definition of $T$,
@@
    In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ for all $t\in\mathbb{N}$. Then $\bar{x}_{t}=x^{\circ}_{t}$, $\bar{y}_{t}=y^{\circ}_{t}$, and $\bar{z}_{t}=z^{\circ}_{t}$ for all $t\in\mathbb{N}$. Also $[\bar{x}_{t}]^{+}=x^{\circ}_{t+1}$, $[\bar{y}_{t}]_{\bar{x}_{t}}^{+}=y^{\circ}_{t+1}$, and $[\bar{z}_{t}]_{\bar{x}_{t}}^{+}=z^{\circ}_{t+1}$ for all $t\in\mathbb{N}$. Thus by definition of $T$,
@@
-C Proof of RR Main Theorem
@@
where the equality between 31 and 32 follows from Corollary 6, the equality between 33 and 34 follows from Lemma 3, and the equality between 34 and 35 follows from Lemma 8. ∎
@@
-D Proof of Optimality Theorem
@@
-E The SRABE and its Policy Gradient
@@
We here closely follow the proof of Theorem 3 in [21], which itself modifies the proofs of the Policy Gradient Theorems in Chapter 13.2 and 13.6 [22]. We only make the minimal modifications required to adapt the PPO algorithm developed previously for the SRBE to on for the SRABE.
@@
where the equality between (37) and (38) comes from rolling out the term $\nabla_{\theta}\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right)$ (see Chapter 13.2 in [22] for details), and where $\mathrm{Pr}(s\to s^{\prime},k,\pi)$ is the probability that under the policy $\pi$, the system is in state $s^{\prime}$ at time $k$ given that it is in state $s$ at time 00. ∎
@@
-F The DO-HJ-PPO Algorithm
In this section, we outline the details of our Actor-Critic algorithm DO-HJ-PPO beyond the details given in Algorithm 1.
-F1 The special Bellman updates and the corresponding GAEs
Akin to previous HJ-RL policy algorithms, namely RCPO [24], RESPO [20] and RCPPO [21], DO-HJ-PPO fundamentally depends on the discounted HJ Bellman updates [18]. To solve the RAA and RR problems with the special rewards defined in Theorems 1 & 2, DO-HJ-PPO utilizes the Reach, Avoid and Reach-Avoid Bellman updates, given by

### v2

@@ 9 Conclusions
$\alpha$   Achievement Speed Results from DOHJ-PPO Experiments........................................................................................................................................................................Achievement Speed Results from DOHJ-PPO Experiments
A   Proof of RAA Main Theorem........................................................................................................................................................................A
B   Proof of RR Main Theorem........................................................................................................................................................................B
C   Proof of Optimality Theorem........................................................................................................................................................................C
D   The SRABE and its Policy Gradient........................................................................................................................................................................D
E   The DOHJ-PPO Algorithm........................................................................................................................................................................E
F   DDQN Demonstration........................................................................................................................................................................F
G   Baselines........................................................................................................................................................................G
H   Details of RAA & RR Experiments: Hopper........................................................................................................................................................................H
I   Details of RAA & RR Experiments: F16........................................................................................................................................................................I
J   Details of RAA & RR Experiments: SafetyGym........................................................................................................................................................................J
K   Details of RAA & RR Experiments: HalfCheetah........................................................................................................................................................................K
L   Contrast with Decompositions in Temporal Logic........................................................................................................................................................................L
M   Broader Impacts........................................................................................................................................................................M
N   Acknowledgments........................................................................................................................................................................N
Achievement Speed Results from DOHJ-PPO Experiments
Here we present additional results for RAA and RR problems solved with DOHJ-PPO. In both settings, DOHJ-PPO out-performs or matches the best of baselines with less tuning and faster arrival. Notably as the difficulty of the problem increases the gap increases significantly with DOHJ-PPO remaining the sole algorithm that can achieve the task in reasonable time and in both RAA and RR categories.
Figure 6: Steps to Success ($\leftarrow$) in RAA and RR Tasks for DOHJ-PPO and Baselines For the same 1000 trajectories in Figure 4, we quantify here the number of steps until achievement of both tasks: reaching without crash afterward in the RAA, reaching both goal in the $R$. DOHJ-PPO is not only competitive but consistently achieves the dual-objective problems in the fewest number of steps.
Proof Notation
@@
We let $\mathbb{N}=\{0,1,\dots\}$ be the set of whole numbers.
We let $\mathbb{{A}}$ be the set of maps from $\mathbb{N}$ to $\mathcal{{A}}$. In other words, $\mathbb{{A}}$ is the set of sequences of actions the agent can choose. Given $\mathbf{a}_{1},\mathbf{a}_{2}\in\mathbb{{A}}$, and $\tau\in\mathbb{N}$, we let $\mathbf{a}_{1},\mathbf{a}_{2}]_{\tau}$ be the element of $\mathbb{{A}}$ for which
$\mathbf{a}_{1},\mathbf{a}_{2}]_{\tau}(t)=\begin{cases}\mathbf{a}_{1}(t)&t<\tau,\\ \mathbf{a}_{2}(t-\tau)&t\geq\tau.\end{cases}$
Similarly, given $a\in\mathcal{{A}}$ and $\mathbf{a}\in\mathbb{{A}}$, we let $a,\mathbf{a}]$ be the element of $\mathbb{{A}}$ for which
$a,\mathbf{a}](t)=\begin{cases}a&t=0,\\ \mathbf{a}(t-1)&t\geq 1.\end{cases}$
@@
The $\cdot,\cdot]_{\tau}$ operation corresponds to concatenating two action sequences (using only the $0^{\text{th}}$ to $(\tau-1)^{\text{st}}$ elements of the first sequence), the $\cdot$$\cdot$] operation corresponds to prepending an action to an action sequence, and the $\cdot|_{\tau}$ operation corresponds to removing the $0^{\text{th}}$ to $(\tau-1)^{\text{st}}$ elements of an action sequence.
@@
Appendix A Proof of RAA Main Theorem
@@
If $y\in S_{n}$, then from (5) and (10), we have that $V^{\pi}(y)=v^{*}(y)=v_{n}^{*}(y)=v_{n+1}^{*}(y)$, and from (7) and (10), we also have that $v_{n}^{*}(y)\geq\alpha_{n}\geq\alpha_{n+1}$. Together these establish (12) when $y\in S_{n}$.
So suppose $y\in S_{n+1}\setminus S_{n}$. First, observe that $v_{n+1}^{*}(y)=\alpha_{n+1}$ by (5). There are now two possibilities. If there is some $a\in\mathcal{{A}}$ for which $f(y,a)\in S_{n}$, then (12) follows from claim 2. If instead, $f(y,a)\notin S_{n}$ for each $a\in\mathcal{{A}}$, then $\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(y,a)\right)=$-$\infty$ by (5) (or if $n=0$ by definition of $v_{0}^{*}$). Thus $\alpha_{n+1}=\min\left\{r_{\textup{RAA}}(y),q(y)\right\}$ by (4), so
@@
We introduce some useful notation we will use throughout the rest of the proof. For each $s\in\mathcal{{S}}$, let $s]^{+}=f(s,\pi(s))$, $y]^{+}_{s}=\max\{y,r\left([s]^{+}\right)\}$, $z]^{+}_{s}=\min\{z,q\left([s]^{+}\right)\}$.
@@
    In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ for $t<T$. Then $\bar{x}_{t}=x^{\circ}_{t}$, $\bar{y}_{t}=y^{\circ}_{t}$, and $\bar{z}_{t}=z^{\circ}_{t}$ for all $t\leq T$. It follows that $\bar{x}_{t}]^{+}=x^{\circ}_{t+1}$, $\bar{y}_{t}]_{\bar{x}_{t}}^{+}=y^{\circ}_{t+1}$, and $\bar{z}_{t}]_{\bar{x}_{t}}^{+}=z^{\circ}_{t+1}$ for all $t\leq T$. Thus by definition of $T$,
@@
    In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ for all $t\in\mathbb{N}$. Then $\bar{x}_{t}=x^{\circ}_{t}$, $\bar{y}_{t}=y^{\circ}_{t}$, and $\bar{z}_{t}=z^{\circ}_{t}$ for all $t\in\mathbb{N}$. Also $\bar{x}_{t}]^{+}=x^{\circ}_{t+1}$, $\bar{y}_{t}]_{\bar{x}_{t}}^{+}=y^{\circ}_{t+1}$, and $\bar{z}_{t}]_{\bar{x}_{t}}^{+}=z^{\circ}_{t+1}$ for all $t\in\mathbb{N}$. Thus by definition of $T$,
@@
A.1 A direct derivation of the RAA Bellman equation
Here, we offer a direct derivation for the RAA Bellman equation. Note, this derivation does not guarantee that the resulting Bellman equation is unique, and is just for intuition for the rigor above.
$v_{\mathrm{RAA}}^{*}(s)$ $=\max_{a_{0},a_{1},\dots}\min\left\{\max_{\tau\in\{0,1,\dots\}}r(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\tau)),\min_{\kappa\in\{0,1,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\}$ (29)
$=\min\left\{q(s),\max_{a_{0},a_{1},\dots}\min\left\{\max_{\tau\in\{0,1,\dots\}}r(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\tau)),\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\}\right\}$
$=\min\left\{q(s),\max_{a_{0},a_{1},\dots}\min\left\{\max\left\{r(s),\max_{\tau\in\{1,2,\dots\}}r(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\tau))\right\},\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\}\right\},$
where $\mathbf{x}_{s}^{a_{0},a_{1},\dots}$ is the system trajectory starting from state $s$ under the sequence of actions $a_{0},a_{1},\dots$. Using the identity $\min\{\max\{a,b\},c\}=\max\{\min\{a,c\},\min\{b,c\}\}$,
$v_{\mathrm{RAA}}^{*}(s)$ $=\min\bigg\{q(s),\max_{a_{0},a_{1},\dots}\max\bigg\{\min\left\{r(s),\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\},$
$\qquad\qquad\qquad\qquad\qquad\qquad\min\left\{\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa)),\max_{\tau\in\{1,2,\dots\}}r(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\tau))\right\}\bigg\}\bigg\}$
$=\min\bigg\{q(s),\max\bigg\{\min\left\{r(s),\max_{a_{0},a_{1},\dots}\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\},$
$\qquad\qquad\qquad\qquad\quad\max_{a_{0},a_{1},\dots}\min\left\{\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa)),\max_{\tau\in\{1,2,\dots\}}r(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\tau))\right\}\bigg\}\bigg\}$
$=\min\bigg\{q(s),\max\bigg\{\min\left\{r(s),\max_{a_{0},a_{1},\dots}\min_{\kappa\in\{1,2,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\},\max_{a}v_{\mathrm{RAA}}^{*}(f(s,a))\bigg\}\bigg\}$
$=\min\bigg\{q(s),\max\bigg\{\min\left\{q(s),r(s),\max_{a_{0},a_{1},\dots}\min_{\kappa\in\{0,1,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\},\max_{a}v_{\mathrm{RAA}}^{*}(f(s,a))\bigg\}\bigg\}$
$=\min\bigg\{q(s),\max\bigg\{\min\left\{r(s),\max_{a_{0},a_{1},\dots}\min_{\kappa\in\{0,1,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))\right\},\max_{a}v_{\mathrm{RAA}}^{*}(f(s,a))\bigg\}\bigg\},$
where in the penultimate step we used the identity $\min\{a,\max\{b,c\}\}=\min\{a,\max\{\min\{a,b\},c\}\}$. Noticing that $v_{\mathrm{A}}^{*}(s)=\max_{a_{0},a_{1},\dots}\min_{\kappa\in\{0,1,\dots\}}q(\mathbf{x}_{s}^{a_{0},a_{1},\dots}(\kappa))$, we have
$v_{\mathrm{RAA}}^{*}(s)$ $=\min\left\{q(s),\max\left\{\min\left\{r(s),v_{\mathrm{A}}^{*}(s)\right\},\max_{a}v_{\mathrm{RAA}}^{*}(f(s,a))\right\}\right\}.$ (30)
This completes the derivation of the RAA Bellman equation.
Lastly, we may note that if define $\tilde{r}(s):=\min\{r(s),v_{\mathrm{A}}^{*}(s)\}$, the above becomes the RA Bellman equation,
$v_{\mathrm{RAA}}^{*}(s)$ $=\min\left\{q(s),\max\left\{\tilde{r}(s),\max_{a}v_{\mathrm{RAA}}^{*}(f(s,a))\right\}\right\}.$ (31)
Appendix B Proof of RR Main Theorem
@@
where the equality between 34 and 35 follows from Corollary 6, the equality between 36 and 37 follows from Lemma 3, and the equality between 37 and 38 follows from Lemma 8. ∎
@@
Appendix C Proof of Optimality Theorem
@@
Appendix D The SRABE and its Policy Gradient
We first justify the SRABE from a theoretical perspective. For each $\gamma\in(0,1)$ and stochastic policy $\pi:\mathcal{{S}}\to\Delta(\mathcal{{A}})$ (with $\Delta(\mathcal{{A}})$ the probability simplex on $\mathcal{{A}}$), let $\tilde{V}_{\mathrm{RA}}^{\gamma,\pi},V_{\mathrm{RA}}^{\gamma,*}:\mathcal{{S}}\to\mathbb{R}$ be the (unique) solutions of the Bellman equations
$\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}(s)=(1-\gamma)\min\left\{r(s),q(s)\right\}+\gamma\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}\left(f(s,a)\right),r(s)\right\},q(s)\right\}\right],$
$V_{\mathrm{RA}}^{\gamma,*}(s)=(1-\gamma)\min\left\{r(s),q(s)\right\}+\gamma\left[\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}V_{\mathrm{RA}}^{\gamma,*}\left(f(s,a)\right),r(s)\right\},q(s)\right\}\right],$
respectively, where $r:\mathcal{{S}}\to\mathbb{R}$ and $q:\mathcal{{S}}\to\mathbb{R}$. Note that the above Bellman equations are indeed $\gamma$-contractive.
Theorem 4.
Let $\gamma\in(0,1)$. Given any $\pi:\mathcal{{S}}\to\Delta(\mathcal{{A}})$, we have
$\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}\leq V_{\mathrm{RA}}^{\gamma,*}.$
Moreover, there exists a $\pi^{*}:\mathcal{{S}}\to\Delta(\mathcal{{A}})$ such that
$\tilde{V}_{\mathrm{RA}}^{\gamma,\pi^{*}}=V_{\mathrm{RA}}^{\gamma,*}.$
Proof.
Let $\pi:\mathcal{{S}}\to\Delta(\mathcal{{A}})$. Define the Bellman operators $B_{\gamma}^{\pi},B_{\gamma}^{*}:\mathbb{R}^{\mathcal{{S}}}\to\mathbb{R}^{\mathcal{{S}}}$ (where $\mathbb{R}^{\mathcal{{S}}}$ is the set of all maps $v:\mathcal{{S}}\to\mathbb{R}$) by
$B_{\gamma}^{\pi}[v](s)=(1-\gamma)\min\left\{r(s),q(s)\right\}+\gamma\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{v\left(f(s,a)\right),r(s)\right\},q(s)\right\}\right],$
$B_{\gamma}^{*}[v](s)=(1-\gamma)\min\left\{r(s),q(s)\right\}+\gamma\left[\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}v\left(f(s,a)\right),r(s)\right\},q(s)\right\}\right],$
respectively. For each $v\in\mathbb{R}^{\mathcal{{S}}}$, we have $B_{\gamma}^{\pi}[v]\leq B_{\gamma}^{*}[v]$. Then $\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}=B_{\gamma}^{\pi}[\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}]\leq B_{\gamma}^{*}[\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}]$ and $V_{\mathrm{RA}}^{\gamma,*}=B_{\gamma}^{*}[V_{\mathrm{RA}}^{\gamma,*}]$.
Since $B_{\gamma}^{*}$ is a contraction and also a monotonic operator ($B_{\gamma}^{*}[v]\leq B_{\gamma}^{*}[w]$ when $v\leq w$), it follows from the comparison principle for Bellman operators that $\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}\leq V_{\mathrm{RA}}^{\gamma,*}$.
Now let $\pi^{*}:\mathcal{{S}}\to\mathcal{{A}}$ be such that $\pi^{*}(s)$ is supported on $\operatorname*{arg\,max}_{a\in\mathcal{{A}}}v(f(s,a))$. Then $\tilde{V}_{\mathrm{RA}}^{\gamma,\pi^{*}}=B_{\gamma}^{\pi^{*}}[\tilde{V}_{\mathrm{RA}}^{\gamma,\pi^{*}}]=B_{\gamma}^{*}[\tilde{V}_{\mathrm{RA}}^{\gamma,\pi^{*}}]$ and $V_{\mathrm{RA}}^{\gamma,*}=B_{\gamma}^{*}[V_{\mathrm{RA}}^{\gamma,*}]$, so that $\tilde{V}_{\mathrm{RA}}^{\gamma,\pi^{*}}=V_{\mathrm{RA}}^{\gamma,*}$. ∎
To understand the significance of the above theorem, recall that when solving RA problems (as is needed during our solution of the RAA problem) we are interested in estimating $V_{\mathrm{RA}}^{\gamma,*}$ in the limit $\gamma\to 1^{-}$ (see Proposition 3 in Fisac et al. [2019]). The above theorem tells us that to obtain $V_{\mathrm{RA}}^{\gamma,*}$ we can search for a (possibly stochastic) policy $\pi$ that maximizes $\tilde{V}_{\mathrm{RA}}^{\gamma,\pi}$. Doing so allows us to use the PPO adaptation described in the DOHJ-PPO algorithm for finding RA value functions. Analogous results hold for the R and A subproblems.
D.1 Policy gradient
By analogy to the SRBE, the SRABE is given by
$\tilde{V}_{\textup{RAA}}^{\pi}(s)=\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right].$ (SRABE)
The corresponding action-value function is
We then have the following proposition:
Proposition 1.
For each $s\in\mathcal{{S}}$ and every $\theta\in\mathbb{R}^{n_{p}}$, we have
with the bracketed term equal to $1$ if the proposition inside is true and 0 otherwise.
Following Hsu et al. [2021], we then define the discounted value and action-value functions with $\gamma\in[0,1)$.
@@ D.1 Policy gradient
We here closely follow the proof of Theorem 3 in So et al. [2024], which itself modifies the proofs of the Policy Gradient Theorems in Chapter 13.2 and 13.6 Sutton and Barto [2018]. We only make the minimal modifications required to adapt the PPO algorithm developed previously for the SRBE to on for the SRABE.
@@
where the equality between (40) and (41) comes from rolling out the term $\nabla_{\theta}\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right)$ (see Chapter 13.2 in Sutton and Barto [2018] for details), and where $\mathrm{Pr}(s\to s^{\prime},k,\pi)$ is the probability that under the policy $\pi$, the system is in state $s^{\prime}$ at time $k$ given that it is in state $s$ at time 0. ∎
@@
Appendix E The DOHJ-PPO Algorithm
In this section, we outline the details of our Actor-Critic algorithm DOHJ-PPO beyond the details given in Algorithm 1.
Algorithm 1 : DOHJ-PPO (Actor-Critic)
1:Composed and Decomposed Actor parameters $\theta$ and $\theta_{i}$, Composed and Decomposed Critic parameters $\omega$ and $\omega_{i}$, GAE $\lambda$, learning rate $\beta_{k}$ and discount factor $\gamma$. Let $B^{\gamma}$ amd $B^{\gamma}_{i}$ represent the Bellman update and decomposed Bellman update for the users choice of problem ($R$ or RAA).
2:Define Composed Actor and Critic $\tilde{Q}$
3:Define Decomposed Actor(s) and Critic(s) $\tilde{Q}_{i}$
4:for $k=0,1,\ldots$ do
5:  for $t=0$ to T1T-1 do
6:   Sample trajectories for $\tau_{t}:\{\hat{s}_{t},a_{t},\hat{s}_{t+1}\}$
7:   Define $\tilde{\ell}(s_{t})$ with Decomposed Critics $\tilde{Q}_{i}(s_{t})$ (Theorems  1 & 2)
8:   Composed Critic update:
$\omega\leftarrow\omega-\beta_{k}\nabla_{\omega}\tilde{Q}(\tau_{t})\cdot\left(\tilde{Q}(\tau_{t})-B^{\gamma}[\tilde{Q},\tilde{r}](\tau_{t})\right)$
9:   Compute Bellman-GAE $A_{HJ}^{\lambda}$ with $B^{\gamma}$
10:   (Standard) update Composed Actor
11:   Decomposed Critic update(s):
12:   Compute Bellman-GAE $A_{i}^{\lambda}$ with $B^{\gamma}_{i}$
13:   (Standard) update Decomposed Actor(s)
14:  end for
15:end for
16:return parameter $\theta$, $\omega$
In Algorithm 1, the Bellman update $B^{\gamma}[\tilde{Q},\tilde{r}$] differs for the RAA task and RR task, and the $B_{i}^{\gamma}[\tilde{Q}$] differs between the reach, avoid, and reach-avoid tasks.
E.1 The special Bellman updates and the corresponding GAEs
Akin to previous HJ-RL policy algorithms, namely RCPO Yu et al. [2022b], RESPO Ganai et al. [2023] and RCPPO So et al. [2024], DOHJ-PPO fundamentally depends on the discounted HJ Bellman updates Fisac et al. [2019]. To solve the RAA and RR problems with the special rewards defined in Theorems 1 & 2, DOHJ-PPO utilizes the Reach, Avoid and Reach-Avoid Bellman updates, given by